OpenStudy (anonymous):

Question 1: Create your own radical expression. Question 2: Convert it to a rational exponent. Question 3: Explain, in complete sentences, how the expression was converted.

OpenStudy (anonymous):

can someone help me please

OpenStudy (anonymous):

i dont know how to create one

OpenStudy (anonymous):

a radical expression is anything of the form\[\sqrt[n]{x}\]where n is an integer and x is a real number. So, pick any numbers for n and x and you have a radical expression. For question 2, we use the fact that \[\sqrt[n]{x}=x^{\frac{1}{n}}\]. So rewirte your created radiacal from question 1 into this form using the same values for n and x.